Circular No. 7878

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN SAGITTARIUS W. Liller, Viña del Mar, Chile, reports his discovery, on two Technical

Pan films (+ orange filter) taken on Apr. 15.354 UT, of an apparent nova (mag 9.2) located at α = 18^{h}00^{m}.1, δ = −30^{o}54′ (equinox 2000.0). A CCD image taken with a 0.25-m reflector by P. Cacella, Brasilia, Brazil, on Apr. 16.093 yields V = 10.1 and α = 17^{h}59^{m}59^{s}.63, δ = −30^{o}53′20′′.5. The closest catalogued object appears to be a very red star in the USNO-A2.0 catalogue that is some 3′′ away (blue mag 19.4, red mag 16.9; position end figures 59^{s}.47, 18′′.7). Additional magnitudes (visual unless otherwise noted): Apr. 11.40, [11.5 (Liller; Tech Pan films); 13.814, [11.0 (K. Kanatsu, Matsue, Shimane, Japan; T-Max 400 film; communicated by T. Kato, Kyoto); 14.78, [8.5 (A. Pearce, Nedlands, W. Australia); 16.129, 9.5 (R. Y. Shida, São Paulo, Brazil); 16.485, 9.8 (M. Linnolt, Honolulu, HI).

SUPERNOVA 2000fu IN MCG –03-38-21 W. D. Li, University of California at Berkeley, reports the discovery

of an apparent supernova on unfiltered CCD archival KAIT images taken during the course of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126). SN 2000fu is located at α = 14^{h}51^{m}12^{s}.67, δ = −20^{o}26′08′′.3 (equinox 2000.0), which is 5′′.0 east and 18′′.1 north of the nucleus of MCG –03-38-21 = ESO 580-G45. Available photometry (all unfiltered): 2000 June 6.3 UT, 16.5; 24.2, 16.8; July 1.2, 16.9. A recent KAIT image taken on 2002 Apr. 13.4 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2002cf IN NGC 4786 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002cf (cf. IAUC 7877), obtained by M. Calkins on Apr. 15.29 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 4786 km/s for the host galaxy, the supernova expansion velocity is ∼ 11 400 km/s for Si II (rest 635.5 nm) and 14 500 km/s for Ca II (rest 395.1 nm). The spectral-feature age of the supernova (cf. IAUC 7774) is 3 ± 2 days before maximum light. The spectrum appears quite similar to that of the subluminous SN 1999by, three days before maximum (see Garnavich et al. 2002, Ap.J., in press,

http://xxx.lanl.gov/abs/astro-ph/?0105490).

2002 April 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green